OpenStudy (sushi121212):

Use expansion by minors to find the value of |P|, in terms of a, b, and c.

OpenStudy (sushi121212):

a b c P= 1 3 5 2 4 6

OpenStudy (jango_in_dtown):

the minor of a will be 3*6-5*4=18-20=-2 that of b will be 1*6-5*2=6-10=-4 that of c will be1*4-2*3=4-6=-2 so the answer is a(-2)-b(-4)+c(-2)=4b-2a-2c

OpenStudy (jango_in_dtown):

@Sushi121212

OpenStudy (sushi121212):

So does that mean that a=-2, b=-4, and c=-2?

OpenStudy (jango_in_dtown):

no |P|=4b-2a-2c

OpenStudy (sushi121212):

Oh, so is the question asking to find P, not a, b, and c? I think I misread the question!! :D

OpenStudy (jango_in_dtown):

yeah

OpenStudy (sushi121212):

Oh, thank you so much!!!